17/Sep/99 Evening shift

Peter Shanahan (shanahan@fnal.gov)
Sat, 18 Sep 1999 00:37:36 -0500

Shift Summary for eve shift, 17/Sep/99

On Shift: J. Hirschauer, P. Shanahan

Hours of beam available at all intensities: 7.5
Hours of beam available at nominal intensity: N/A
Hours of E799 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 7.0

Accounting of lost time: kfrend for crate 53 every 20-30', intitializing runs

========================================================================
Runs ended this shift:
Run 14580 - - BEAM_MUON - DC/TRD plateau
Run 14581 - 440K - BEAM_799 - First E799 hadron beam
Run 14582 - 2.1M - BEAM_799 - Rate vs. absorber studies
Run 14586 - 1.2M - BEAM_799 - 0 absorber
Run 14588 - 0.6M - BEAM_799 - 0 absorber, new prescales/crate map
Run 14589 - 1.0M - BEAM_799 - 10'' Be Absorber

Tapes brought to FCC this shift: None
(including tapes for online split)

========================================================================
Accesses during shift: None

========================================================================
Overview of shift:
DC/TRD muon plateau and E799 hadron commissioning. Studies of 0, 10, and
20'' Be absorber at 7E12. Data is consistent so far with 97 experience.
Work in progress...

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
16:30 DC/TRD muon studies finished. Set NM4AN to +600 A,
request sweepers set for E799 hadron mode
14581 0 16:40 Run control still expects Pt kick of 205 MeV.
DIS value is 153.29. We continue.
16:45 Request 2E12. Jaws are set to 2cm x 2cm. 20'' Be
absorber in. Beamstop out.
14581 1 17:00 Start Beam_799 run.
17:10 3 new Ar/Ethane bottles online on west side
14581 13 17:15 Jaw collimator opened to 4cm x 4cm.
14581 21 Jaw collimator opened to 11cm x 11cm (full size)
14581 48 18:01 CA HV turned on (BT)
14581 18:04 Request 4E12
14581 18:07 ET2, ET3 tuned for 18, 25 GeV (JCH)
14581 18:30 End run. Change Ke3 and accid prescales
18:40 Request 7E12
14582 1 18:45 Begin run @ 7E12, 20'' of Be absorber.
14582 22 19:13 Change Be absorber to 10''.
19:17 E/p has been 1% low. Is this error in PT kick?
14582 19:45 End run.
19:55 ET2 changed to 16 GeV for use in B10.
20:40 Attempt to initialize BEAM_799 w/onsplit. Many
errors -> back to non-onsplit configuration. (AT)
14586 20:45 Start run with NO Be absorber. 7E12
14588 1 21:25 New crate map loaded/new Fera sprs. to reduce data vol.
+prescale B10 by 2. Still ~3 sec deadtime/spill.
14589 1 22:12 New run with 10'' Be absorber.
14590 1 23:10 New run with 20'' Be absorber.


========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift

o Page Arun if down for ~>1 hour

o Access from outside .fnal.gov to ktev3 seems restricted. This does not
appear to be a network problem. Probably should NOT call helpdesk
about it overnight...


General issues for near term:

o Accidental counter: little or no RF structure visible. Move 90^o monitor?
o STT: is it working?
o Data volume problems at 7E12, with ABS=0.
o OLS destination problems
o E/p low by 1%. (Hints of low PT kick +-0 dalitz as well)
o Tape writing is not balanced. Work in progress.